OpenStudy (anonymous):

You have one type of chocolate that sells for $2.40/lb and another type of chocolate that sells for $4.00/lb. You would like to have 4.8 lbs of a chocolate mixture that sells for $3.40/lb. How much of each chocolate will you need to obtain the desired mixture?

OpenStudy (shinalcantara):

let x=the amount of chocolate @ 2.4 y=the amount of chocolate @ 4 2.4x+4y=4.8(3.40) ----eqn 1 x+y=4.8 -----eqn 2 x=4.8-y -----eqn 3 substitute eqn 3 to eqn 1 2.4(4.8-y)+4y=16.32 11.52-2.4y+4y=16.32 1.6y=4.8 y=3 lbs substitute to eqn 3 x=4.8-3 x=1.8

OpenStudy (anonymous):

you rock!!!:)

OpenStudy (anonymous):

for some odd reason my review answer key says the answer is: 3.675 lbs. at $2.40/pound and 1.125 lbs. at $4.00/pound. Does this mean they are wrong ? lol

OpenStudy (shinalcantara):

there could have been a mistake on the answer key for as long as your given problem is concerned.. ;-)